(x^2+y^2)y'=y^2


1
Expert's answer
2021-02-16T13:46:53-0500

Solution.

"(x^2+y^2)y'=y^2."


Make a replacement

"u(x)=\frac{y(x)}{x}," from here "y(x)=xu(x)," and "y'=u(x)+xu'(x)."

Substitute the obtained expressions into the equation:

"(x^2+x^2u^2)(u+xu')=x^2u^2,"

"x^2u+x^2u^3+x^3u'+x^3u^2u'-x^2u^2=0,"

"x^3u'(1+u^2)+x^2(u^3-u^2+u)=0."

Divide both parts of the equation by "u^3-u^2+u" and write the equation in the form:


"\frac{(u^2+1)u'}{u^3-u^2+u}=-\frac{1}{x}."

Thus we obtain a separable differential equation. Let's solve it:


"\frac{dx(u^2+1)\frac{du}{dx}}{u^3-u^2+u}=-\frac{dx}{x},""\frac{du(u^2+1)}{u^3-u^2+u}=-\frac{dx}{x},""\int\frac{u^2+1}{u^3-u^2+u}du=\int-\frac{1}{x}dx,""\int\frac{1}{u^2-u+1}du+\int\frac{1}{u}du=\int-\frac{1}{x}dx,""\frac{2\arctan{\frac{2u-1}{\sqrt{3}}}}{\sqrt{3}}+\ln{|u|}=-\ln{|x|}+C."

where C is a constant.

Returning to the replacement we obtain the solution of the given equation:


"\frac{2\arctan{\frac{2\frac{y(x)}{x}-1}{\sqrt{3}}}}{\sqrt{3}}+\ln(|\frac{y(x)}{x}|)=-\ln{|x|}+C,"

or


"\frac{2\sqrt{3}\arctan{\frac{\sqrt{3}(2\frac{y(x)}{x}-1)}{3}}}{3}+\ln(|y(x)|)=C."

Answer.

"\frac{2\sqrt{3}\arctan{\frac{\sqrt{3}(2\frac{y(x)}{x}-1)}{3}}}{3}+\ln(|y(x)|)=C."